how would you solve an equation with variables on both sides for example
12-2a=-5a-9

Answer by wuwei96815(245)

You cannot solve this equation until you get all the variables on the same side.

12-2a = -5a-9 First, add 2a to each side of the equation.

You get: 12-2a+2a = -5a-9+2a
Simplified it becomes: 12 = -3a-9 Now, add 9 to each side.
You get: 12+9 = -3a-9+9
Simplified it becomes: 21 = -3a
21/-3 = a
-7 = a
You can check the correctness of this answer by substituting -7 for a in the original equation.